# The Three Regimes of Offline-to-Online Reinforcement Learning

### v2

@@ Abstract
Offline-to-online reinforcement learning (RL) has emerged as a practical paradigm that leverages offline datasets for pretraining and online interactions for fine-tuning. However, its empirical behavior is highly inconsistent: design choices of online-fine tuning that work well in one setting can fail completely in another. We propose a stability–plasticity principle that can explain this inconsistency: we should preserve the knowledge of pretrained policy or offline dataset during online fine-tuning, whichever is better, while maintaining sufficient plasticity. This perspective identifies three regimes of online fine-tuning, each requiring distinct stability properties. We validate this framework through a large-scale empirical study, finding that the results strongly align with its predictions in 45 of 63 cases. This work provides a principled framework for guiding design choices in offline-to-online RL based on the relative performance of the offline dataset and the pretrained policy.
@@ 1 Introduction
Reinforcement learning (RL) has achieved impressive successes in a variety of domains (mnih2015human; AlphaGo_Zero; degrave2022magnetic), but its reliance on large amounts of online interaction often makes direct application to real-world problems challenging. To address this challenge, recent research has turned to leveraging pre-collected datasets for offline learning through offline RL (levine2020offline) or imitation learning (osa2018algorithmic). This paradigm reduces the need for costly or unsafe online exploration by providing a strong initial policy trained entirely from offline data. However, policies trained purely offline are often suboptimal and fail to generalize to states outside the dataset’s support, making online fine-tuning essential. Offline-to-online RL addresses this issue by first pretraining an agent on an offline dataset and then fine-tuning it with additional online interactions to further improve performance.
Figure 1: Comparison between WSRL (pretrained policy only) and RLPD (offline dataset only) on two representative offline-to-online RL tasks. All learning curves are shown as mean $\pm$ 95% CI.
While offline-to-online RL has led to promising results (nair2020awac; lee2022offline), online RL fine-tuning suffers from highly inconsistent empirical behavior: design choices that work well in one setting can fail completely in another. For example, as shown in Figure 1, on D4RL tasks (fu2020d4rl) such as antmaze-large-play-v2, Warm-Start RL (WSRL) (zhou2024efficient), which relies on the pretrained policy and discards the offline dataset during online fine-tuning, substantially outperforms RL with Prior Data (RLPD) (ball2023efficient), which uses only the offline dataset. In contrast, on D4RL tasks such as relocate-binary-v0, the opposite pattern emerges, with RLPD outperforming WSRL by a wide margin. These seemingly inconsistent outcomes raise one fundamental question:
@@
To answer this question, we propose a stability–plasticity principle for offline-to-online RL that can explain these seemingly inconsistent outcomes, which is inspired by prior research on plasticity and stability in neuroscience (mcclelland1995there) and machine learning  (kirkpatrick2017overcoming; wolczyk2024fine; dohare2024loss). Guided by the principle, effective fine-tuning requires a careful balance between stability and plasticity. Stability refers to the preservation of useful prior knowledge, ensuring that competencies acquired during pretraining are not substantially degraded. Plasticity, in contrast, denotes the capacity of the model to adapt flexibly and efficiently to new data. Furthermore, we identify two distinct forms of stability in offline-to-online RL: stability around the pretrained policy $\pi_{0}$, which emphasizes preserving knowledge explicitly encoded in the policy parameters, and stability around the offline dataset $\mathcal{D}$, which emphasizes retaining knowledge implicitly encoded in offline data. As stability and plasticity are inherently in trade-off, this distinction indicates that fine-tuning is more efficient when stability is enhanced with respect to the stronger source of prior knowledge, whether it is the pretrained policy or the offline dataset.
@@
  • We conduct an extensive empirical study across 63 settings to validate and analyze this framework.
@@ 2 Preliminary: Offline-to-Online RL
The offline dataset $\mathcal{D}$ may be reused (nakamoto2023cal) or discarded (zhou2024efficient) during fine-tuning. The objective of offline-to-online RL is to co-design $(A_{\text{off}},A_{\text{on}})$ such that the final policy $\pi_{N}$ maximizes $J(\pi_{N})$. In this work, we focus on understanding and improving the online RL fine-tuning component $A_{\text{on}}$ by fixing the offline pretraining component $A_{\text{off}}$.
@@ 3 A Stability–Plasticity Principle on Knowledge Acquisition
This section introduces a formal framework for reasoning about fine-tuning in offline-to-online RL. Our goal is to characterize when and how online training leads to improvements over the offline initialization or degradations of what was already learned. We define two complementary properties of online fine-tuning: stability, the ability to preserve previously acquired performance, and plasticity, the capacity to improve further. These are grounded in the notion of a knowledge level, understood as the expected return encoded either in the dataset or in the pretrained policy. We show that final performance admits a decomposition into three terms — prior knowledge, stability, and plasticity. We leverage this perspective provides both theoretical insight and practical guidance.
@@ 3.1 Quantifying Prior Knowledge
where $L$ is the number of trajectories in $\mathcal{D}$, and $r_{k,t}$ is the reward at time step $t$ in the $k$-th trajectory. This measure provides a scalar summary of the return level encoded in the dataset. [^1] We do not assume access to additional quantities such as data coverage or performance surrogates (e.g., dense rewards for sparse-reward tasks), which could further enrich the quantification of prior knowledge.
Knowledge from the pretrained policy: $J(\pi_{0})$. Let $\pi_{0}$ be the policy produced by applying an offline RL algorithm $A_{\text{off}}$ to $\mathcal{D}$. Its performance, $J(\pi_{0})$, reflects not only the quality of the dataset but also the inductive biases of $A_{\text{off}}$ and the difficulty of the MDP $\mathcal{M}$. This policy can provide a strong initialization for online fine-tuning, though it does not necessarily dominate the dataset baseline. We therefore consider both $J(\pi_{\mathcal{D}})$ and $J(\pi_{0})$ jointly as candidate sources of prior knowledge.
3.2 Knowledge Decomposition and the Three Regimes of Offline-to-Online RL
@@
$\vskip-10.00002pt\text{Stability}(l)\vcentcolon=\min\left(\min_{0\leq n\leq N}J(\pi_{n})-l,0\right).$ (2)
@@
In our setting, the appropriate reference level is the best performance available from the offline pretraining phase, either from the dataset or from the pretrained policy: {myframe}
@@
$\underbrace{\max_{0\leq n\leq N}J(\pi_{n})}_{\text{Final knowledge}}=\underbrace{J^{*}_{\text{off}}}_{\text{Prior knowledge}}+\underbrace{\text{Stability}(J^{*}_{\text{off}})}_{\text{Degradation on prior knowledge}\leq 0}+\underbrace{\text{Plasticity}}_{\text{Online knowledge}\geq 0}.$ (6)
@@
This equation states that the final knowledge an agent acquires after offline pretraining and online fine-tuning is the outcome of three interacting components. (1) The first term is the prior knowledge provided by the offline phase, either through the dataset or the pretrained policy, whichever is stronger. (2) The second term measures stability, which records whether this prior knowledge is preserved or degraded during fine-tuning; it is always non-positive since performance can at best be maintained but not exceeded by this term. (3) The third term captures plasticity, the additional knowledge acquired through online interaction, which is non-negative by definition. Therefore, the improvement over the prior knowledge is given by the sum of plasticity and stability.
@@
The three regimes of offline-to-online RL. With prior knowledge fixed, the objective of online fine-tuning is to improve final performance by maintaining stability with respect to $\max(J(\pi_{0}),J(\pi_{\mathcal{D}}))$ while ensuring sufficient plasticity. Based on this perspective, we identify three regimes for the online fine-tuning phase of offline-to-online RL: Superior: where $J(\pi_{0})>J(\pi_{\mathcal{D}})$; Comparable: where $J(\pi_{0})\approx J(\pi_{\mathcal{D}})$; Inferior: where $J(\pi_{0})<J(\pi_{\mathcal{D}})$. These regimes are intended to reflect substantial differences in $J(\pi_{0})$ and $J(\pi_{\mathcal{D}})$, since small performance gaps may not be meaningful and therefore should not determine regime assignment. We additionally explore alternative taxonomies using different metrics; the full details are given in Appendix C.
@@
This regime taxonomy provides a principled framework for reasoning about the stability–plasticity trade-off in offline-to-online RL. It clarifies which source of prior knowledge should anchor stability in a given setting, as shown in Figure 2. In the Superior Regime, stability relative to $\pi_{0}$ should be prioritized, because $\pi_{0}$ contains more useful knowledge than $\mathcal{D}$. In the Inferior Regime, stability relative to $\mathcal{D}$ should be emphasized, as $\mathcal{D}$ contains more useful knowledge than $\pi_{0}$. In the Comparable Regime, both baselines provide similar knowledge, so preserving either helps. At the same time, maintaining sufficient plasticity across all regimes is essential for efficient fine-tuning. By categorizing design choices according to the stability or plasticity they promote, our framework turns what previously appeared as a disconnected set of practices into a structured and predictable landscape, providing a principled foundation for future algorithm design.
4 Design Choices in Improving Stability or Plasticity
Building on the stability–plasticity principle and the regime taxonomy, we now examine concrete design choices that instantiate these principles. To study these factors systematically, we isolate and analyze representative modules, each of which targets one of the three directions: stability around $\pi_{0}$, stability around $\mathcal{D}$, or increased plasticity. Although many existing offline-to-online RL algorithms combine multiple components that promote these aspects simultaneously, such entanglement makes it difficult to attribute effects to a specific source. To obtain a clearer scientific understanding, we isolate and analyze representative modules individually.
@@ 4 Design Choices in Improving Stability or Plasticity
Minimal baseline. We begin with defining a naive online RL fine-tuning baseline that serves as the reference point for introducing additional components, which is intentionally minimalist. It applies a standard online RL algorithm (e.g., SAC (haarnoja2018soft)) initialized with an offline-pretrained agent, without any further modifications. This baseline anchors the analysis and makes the marginal effect of each added component interpretable.
@@ 4.1 Stability Relative to the Offline Dataset $\mathcal{D}$
A common strategy for incorporating offline data during fine-tuning is to reuse the offline dataset together with newly collected online transitions. One approach initializes the replay buffer with the entire offline dataset, after which new online experiences are appended as the agent interacts with the environment. In this case, the ratio of offline to online data is determined by the dataset size and gradually shifts toward online data as training progresses. An alternative approach maintains two separate replay buffers: one fixed buffer containing the offline dataset and another buffer for online experiences. During training, each batch is sampled from both buffers according to a specified offline data ratio $\alpha$. For instance, RLPD (ball2023efficient) uses $\alpha=0.5$, corresponding to a symmetric 50% offline and 50% online sampling ratio.
@@ 4.2 Stability with Respect to the Pretrained Policy $\pi_{0}$
Online data warmup. Before applying gradient updates, the agent $\pi_{0}$ first collects a larger amount of online data ($K$ steps). This strategy reduces the mismatch between the pretraining distribution and the online data distribution, lowering the chance that early updates overwrite prior knowledge. It was introduced in WSRL (zhou2024efficient).
Offline RL regularization. Fine-tuning can also reuse the same offline RL algorithm that produced the pretrained policy, thereby inheriting its conservative regularization. This regularization penalizes state-action pairs outside the online data. Since the online data is collected by the sequence of policies from $\pi_{0}$ to $\pi_{N}$, the regularization implicitly anchors learning around the region visited by $\pi_{0}$, even if $\pi_{0}$ is not stored during fine-tuning. When offline data $\mathcal{D}$ is also used with regularization, we consider it as promoting stability towards both $\pi_{0}$ and $\mathcal{D}$; since $\pi_{0}$ is derived from $\mathcal{D}$, this setting tends to have the strongest stability. Such regularization is widely adopted in prior work (nair2020awac; kostrikov2021offline; tarasov2023revisiting; nakamoto2023cal).
@@ 4.3 Plasticity: Parameter Reset
A direct way to increase plasticity is to periodically reset network parameters while retaining the accumulated training data. Randomly initialized networks tend to exhibit higher plasticity than pretrained ones (nikishin2022primacy), although this strategy may come at the cost of severe forgetting and performance degradation. In the context of offline-to-online RL, parameter reset can be interpreted as starting from any pretrained agent $\pi_{0}$ and then resetting its weights to a randomly initialized agent $\pi_{1}$. While this approach severely degrades initial performance (i.e., $J(\pi_{1})-J(\pi_{0})$ is highly negative), it can significantly enhances plasticity. RLPD (ball2023efficient) directly trains an online RL agent from random initialization, which can be viewed as pretraining followed by fully parameter reset before fine-tuning.
@@ 5 Empirical Study
We study a diverse set of benchmark tasks and pretraining dataset compositions, following the experimental protocols of prior work (nakamoto2023cal; zhou2024efficient). Specifically, we include MuJoCo locomotion, AntMaze, Kitchen, and Adroit domains from D4RL (fu2020d4rl), covering a total of 21 dataset-task compositions. All experiments are conducted with 10 random seeds to ensure statistical reliability.
Offline pretraining phase. We employ two representative offline RL algorithms, Calibrated Q-Learning (CalQL) (nakamoto2023cal) and ReBRAC (tarasov2023revisiting), as well as behavior cloning (BC) (schaal1996learning) using a deterministic policy. To pair a behavior-cloned policy with a critic, we pretrain the critic by Fitted Q Evaluation (FQE) (voloshin2019empirical) after BC. Combining the 21 dataset-task compositions with these 3 pretraining algorithms yields 63 experimental settings in total. Each setting is defined by a specific combination of pretraining algorithm, dataset, and task. In this work, since we focus on cases where the offline dataset and the task are from the same MDP, the pretraining algorithm and dataset are sufficient to uniquely specify a setting.
@@
Online fine-tuning phase. To ensure consistency between offline pretraining phase and online fine-tuning phase, we fine-tune each agent using the corresponding base algorithm. Specifically, we fine-tune CalQL-pretrained agents using SAC (haarnoja2018soft) and ReBRAC-pretrained agents using TD3 (fujimoto2018addressing). For the deterministic BC pretraining, we use TD3 for fine-tuning to match its deterministic actor structure, and use ReBRAC when applying regularization.
Since evaluating every possible design and their combinations is infeasible, and we have grouped them into four categories: the minimal baseline, $\pi_{0}$-centric methods, $\mathcal{D}$-centric methods, and mixed $\pi_{0}+\mathcal{D}$ methods. We then evaluate six representative methods spanning these four categories, which collectively capture the key design choices explored in prior offline-to-online RL literature.
@@
  • Mixed $\bm{\pi_{0}+\mathcal{D}}$ methods: The baseline combined with offline data replay and offline RL regularization, a combination widely adopted in prior work (nair2020awac; kostrikov2021offline; tarasov2023revisiting; nakamoto2023cal).

### v3

@@ Abstract
Offline-to-online reinforcement learning (RL) has emerged as a practical paradigm that leverages offline datasets for pretraining and online interactions for fine-tuning. However, its empirical behavior is highly inconsistent: design choices of online fine-tuning that work well in one setting can fail completely in another. We propose a stability–plasticity principle that can explain this inconsistency: we should preserve the knowledge of pretrained policy or offline dataset during online fine-tuning, whichever is better, while maintaining sufficient plasticity. This perspective identifies three regimes of online fine-tuning, each requiring distinct stability properties. We validate this framework through a large-scale empirical study, finding that the results strongly align with its predictions in 45 of 63 cases, with only 3 opposite mismatches. This work provides a principled framework for guiding design choices in offline-to-online RL based on the relative performance of the offline dataset and the pretrained policy.
Machine Learning, ICML
@@ 1 Introduction
Reinforcement learning (RL) has achieved impressive successes in a variety of domains (Mnih et al., 2015; Silver et al., 2017; Degrave et al., 2022), but its reliance on large amounts of online interaction often makes direct application to real-world problems challenging. To address this challenge, recent research has turned to leveraging pre-collected datasets for offline learning through offline RL (Levine et al., 2020) or imitation learning (Osa et al., 2018). This paradigm reduces the need for costly or unsafe online exploration by providing a strong initial policy trained entirely from offline data. However, policies trained purely offline are often suboptimal and fail to generalize to states outside the dataset’s support, making online fine-tuning essential. Offline-to-online RL addresses this issue by first pretraining an agent on an offline dataset and then fine-tuning it with additional online interactions to further improve performance.
Figure 1: Comparison between WSRL (pretrained policy only) and RLPD (offline dataset only) on two representative offline-to-online RL tasks. Learning curves are shown as mean $\pm$ 95% CI.
While offline-to-online RL has led to promising results (Nair et al., 2020; Lee et al., 2022), online RL fine-tuning suffers from highly inconsistent empirical behavior: design choices that work well in one setting can fail completely in another. For example, as shown in Figure 1, on D4RL tasks (Fu et al., 2020) such as antmaze-large-play-v2, Warm-Start RL (WSRL) (Zhou et al., 2024), which relies on the pretrained policy and discards the offline dataset during online fine-tuning, substantially outperforms RL with Prior Data (RLPD) (Ball et al., 2023), which uses only the offline dataset. In contrast, on D4RL tasks such as relocate-binary-v0, the opposite pattern emerges, with RLPD outperforming WSRL by a wide margin. These seemingly inconsistent outcomes raise one fundamental question:
@@
To answer this question, we propose a stability–plasticity principle for offline-to-online RL that can explain these seemingly inconsistent outcomes, which is inspired by prior research on plasticity and stability in neuroscience (McClelland et al., 1995) and machine learning  (Kirkpatrick et al., 2017; Wolczyk et al., 2024; Dohare et al., 2024). Guided by the principle, effective fine-tuning requires a careful balance between stability and plasticity. Stability refers to the preservation of useful prior knowledge, ensuring that competencies acquired during pretraining are not substantially degraded. Plasticity, in contrast, denotes the capacity of the model to adapt flexibly and efficiently to new data. Furthermore, we identify two distinct forms of stability in offline-to-online RL: stability around the pretrained policy $\pi_{0}$, which emphasizes preserving knowledge explicitly encoded in the policy parameters, and stability around the offline dataset $\mathcal{D}$, which emphasizes retaining knowledge implicitly encoded in offline data. As stability and plasticity are inherently in trade-off, this distinction indicates that fine-tuning is more efficient when stability is enhanced with respect to the stronger source of prior knowledge, whether it is the pretrained policy or the offline dataset.
@@
  • We conduct an extensive empirical study across 63 settings to validate this framework. We also show that our raw-return-based taxonomy consistently outperforms alternatives based on other metrics.
@@ 2 Preliminary: Offline-to-Online RL
The offline dataset $\mathcal{D}$ may be reused (Nakamoto et al., 2023) or discarded (Zhou et al., 2024) during fine-tuning. The objective of offline-to-online RL is to co-design $(A_{\text{off}},A_{\text{on}})$ such that the final policy $\pi_{N}$ maximizes $J(\pi_{N})$. In this work, we focus on understanding and improving the online RL fine-tuning component $A_{\text{on}}$ by fixing the offline pretraining component $A_{\text{off}}$.
@@ 3 A Stability–Plasticity Principle on Knowledge Acquisition
This section introduces a formal framework for reasoning about fine-tuning in offline-to-online RL. Our goal is to characterize when and how online training leads to improvements over the offline initialization or degradations of what was already learned. We define two complementary properties of online fine-tuning: stability, the ability to preserve previously acquired performance, and plasticity, the capacity to improve further. These are grounded in the notion of a knowledge level, understood as the expected return encoded either in the dataset or in the pretrained policy. We show that final performance admits a decomposition into three terms — prior knowledge, stability, and plasticity. This perspective provides both theoretical insight and practical guidance.
@@ 3.1 Quantifying Prior Knowledge
where $L$ is the number of trajectories in $\mathcal{D}$, and $r_{k,t}$ is the reward at time step $t$ in the $k$-th trajectory. This measure provides a scalar summary of the return encoded in the dataset. In Section 5.5, we compare this simple metric with alternative regime taxonomies, including dense reward proxies.
Knowledge from the pretrained policy: $J(\pi_{0})$. The performance $J(\pi_{0})$ reflects not only the quality of the dataset but also the inductive biases of the offline RL algorithm $A_{\text{off}}$ and the difficulty of the MDP $\mathcal{M}$. This policy can provide a strong initialization for online fine-tuning, though it does not necessarily dominate the dataset baseline. We therefore consider both $J(\pi_{\mathcal{D}})$ and $J(\pi_{0})$ jointly as candidate sources of prior knowledge.
3.2 Knowledge Decomposition and Three Regimes
@@
$\text{Stability}(l)\vcentcolon=\min\left(\min_{0\leq n\leq N}J(\pi_{n})-l,0\right).$ (2)
@@
In our setting, the appropriate reference level is the best performance available from the offline pretraining phase, either from the dataset or from the pretrained policy:
@@
$\underbrace{\max_{0\leq n\leq N}J(\pi_{n})}_{\text{Final knowledge}}=\underbrace{J^{*}_{\text{off}}}_{\text{Prior}}+\underbrace{\text{Stability}(J^{*}_{\text{off}})}_{\text{Degradation}\leq 0}+\underbrace{\text{Plasticity}}_{\text{ knowledge}\geq 0}.$
@@
This equation states that the final knowledge an agent acquires is the outcome of three interacting components. (1) The first term is the prior knowledge provided by the offline phase, either through the dataset or the pretrained policy, whichever is stronger. (2) The second term measures stability, which records whether this prior knowledge is preserved or degraded during fine-tuning; it is always non-positive since performance can at best be maintained but not exceeded by this term. (3) The third term captures plasticity, the additional knowledge acquired through online interaction, which is non-negative by definition. Therefore, the improvement over the prior knowledge is given by the sum of plasticity and stability.
@@
The three regimes of offline-to-online RL. With prior knowledge fixed, the objective of online fine-tuning is to improve final performance by balancing between stability with respect to $\max(J(\pi_{0}),J(\pi_{\mathcal{D}}))$ and sufficient plasticity. Based on this perspective, we identify three regimes for the online fine-tuning phase: Superior: where $J(\pi_{0})>J(\pi_{\mathcal{D}})$; Comparable: where $J(\pi_{0})\approx J(\pi_{\mathcal{D}})$; Inferior: where $J(\pi_{0})<J(\pi_{\mathcal{D}})$. These regimes are intended to reflect substantial differences in $J(\pi_{0})$ and $J(\pi_{\mathcal{D}})$, since small performance gaps may not be meaningful and therefore should not determine regime assignment.
@@
This regime taxonomy provides a principled framework for reasoning about the stability–plasticity trade-off in offline-to-online RL. It clarifies which source of prior knowledge should anchor stability in a given setting, as shown in Figure 2. In the Superior Regime, stability relative to $\pi_{0}$ should be prioritized, because $\pi_{0}$ contains more useful knowledge than $\mathcal{D}$. In the Inferior Regime, stability relative to $\mathcal{D}$ should be emphasized, as $\mathcal{D}$ contains more useful knowledge than $\pi_{0}$. In the Comparable Regime, both baselines provide similar knowledge, so preserving either helps. At the same time, maintaining sufficient plasticity across all regimes is essential for efficient fine-tuning.
4 Design Choices in Stability and Plasticity
Building on the stability–plasticity principle and the regime taxonomy, we analyze concrete design choices for online RL fine-tuning. By categorizing methods according to whether they promote stability around the pretrained policy $\pi_{0}$, stability around the offline dataset $\mathcal{D}$, or increased plasticity, our framework organizes previously disparate practices into a structured landscape. Because many existing algorithms entangle these components, we isolate and analyze representative modules individually to clarify their distinct effects.
@@ 4 Design Choices in Stability and Plasticity
Minimal baseline. We begin with defining a naive online RL fine-tuning baseline that serves as the reference point for introducing additional components, which is intentionally minimalist. It applies a standard online RL algorithm (e.g., SAC (Haarnoja et al., 2018)) initialized with an offline-pretrained agent, without any further modifications. This baseline anchors the analysis and makes the marginal effect of each added component interpretable.
@@ 4.1 Stability Relative to the Offline Dataset $\mathcal{D}$
A common strategy for incorporating offline data during fine-tuning is to reuse the offline dataset together with newly collected online transitions. One approach initializes the replay buffer with the entire offline dataset, after which new online experiences are appended as the agent interacts with the environment. In this case, the ratio of offline to online data is determined by the dataset size and gradually shifts toward online data as training progresses. An alternative approach maintains two separate replay buffers: one fixed buffer containing the offline dataset and another buffer for online experiences. During training, each batch is sampled from both buffers according to a specified offline data ratio $\alpha$. For instance, RLPD (Ball et al., 2023) uses $\alpha=0.5$, corresponding to a symmetric 50% offline and 50% online sampling ratio.
@@ 4.2 Stability with Respect to the Pretrained Policy $\pi_{0}$
Online data warmup. Before applying gradient updates, the agent $\pi_{0}$ first collects a larger amount of online data ($K$ steps). This strategy reduces the mismatch between the pretraining distribution and the online data distribution, lowering the chance that early updates overwrite prior knowledge. It was introduced in WSRL (Zhou et al., 2024).
Offline RL regularization. Fine-tuning can also reuse the same offline RL algorithm that produced the pretrained policy, thereby inheriting its conservative regularization. This regularization penalizes state-action pairs outside the online data. Since the online data is collected by the sequence of policies from $\pi_{0}$ to $\pi_{N}$, the regularization implicitly anchors learning around the region visited by $\pi_{0}$, even if $\pi_{0}$ is not stored during fine-tuning. When offline data $\mathcal{D}$ is also used with regularization, we consider it as promoting stability towards both $\pi_{0}$ and $\mathcal{D}$; since $\pi_{0}$ is derived from $\mathcal{D}$, this setting tends to have the strongest stability. Such regularization is widely adopted in prior work (Nair et al., 2020; Kostrikov et al., 2021; Tarasov et al., 2023; Nakamoto et al., 2023).
@@ 4.3 Plasticity: Parameter Reset
A direct way to increase plasticity is to periodically reset network parameters while retaining the accumulated training data. Randomly initialized networks tend to exhibit higher plasticity than pretrained ones (Nikishin et al., 2022), although this strategy may come at the cost of severe forgetting and performance degradation. In the context of offline-to-online RL, parameter reset can be interpreted as starting from any pretrained agent $\pi_{0}$ and then resetting its weights to a randomly initialized agent $\pi_{1}$. While this approach severely degrades initial performance (i.e., $J(\pi_{1})-J(\pi_{0})$ is highly negative), it significantly enhance plasticity. RLPD (Ball et al., 2023) directly trains an online RL agent from random initialization, which can be viewed as pretraining followed by full parameter reset before fine-tuning.
@@ 5 Empirical Study
We study a diverse set of benchmark tasks and pretraining dataset compositions, following the experimental protocols of prior work (Nakamoto et al., 2023; Zhou et al., 2024). Specifically, we include MuJoCo locomotion, AntMaze, Kitchen, and Adroit domains from D4RL (Fu et al., 2020), covering a total of 21 dataset-task compositions. All experiments are conducted with 10 random seeds to ensure statistical reliability.
Offline pretraining phase. We employ two representative offline RL algorithms, Calibrated Q-Learning (CalQL) (Nakamoto et al., 2023) and ReBRAC (Tarasov et al., 2023), as well as behavior cloning (BC) (Schaal, 1996) using a deterministic policy. To pair a behavior-cloned policy with a critic, we pretrain the critic by Fitted Q Evaluation (FQE)  (Le et al., 2019) after BC. Combining the 21 dataset-task compositions with these 3 pretraining algorithms yields 63 experimental settings in total. Each setting is defined by a specific combination of pretraining algorithm, dataset, and task. In this work, since we focus on cases where the offline dataset and the task are from the same MDP, the pretraining algorithm and dataset are sufficient to uniquely specify a setting.
@@
Online fine-tuning phase. To ensure consistency between offline pretraining phase and online fine-tuning phase, we fine-tune each agent using the corresponding base algorithm. Specifically, we fine-tune CalQL-pretrained agents using SAC (Haarnoja et al., 2018) and ReBRAC-pretrained agents using TD3 (Fujimoto et al., 2018). For the deterministic BC pretraining, we use TD3 for fine-tuning to match its deterministic actor structure, and use ReBRAC when applying regularization.
Since evaluating every possible design and their combinations is infeasible, we have grouped them into four categories: the minimal baseline, $\pi_{0}$-centric methods, $\mathcal{D}$-centric methods, and mixed $\pi_{0}+\mathcal{D}$ methods. We then evaluate six representative methods spanning these four categories, which collectively capture the key design choices explored in prior offline-to-online RL literature.
@@
  • Mixed $\bm{\pi_{0}+\mathcal{D}}$ methods: The baseline combined with offline data replay and offline RL regularization, a combination widely adopted in prior work (Nair et al., 2020; Kostrikov et al., 2021; Tarasov et al., 2023; Nakamoto et al., 2023).